# Another look at inference after prediction

### v5

@@ 3.2 Asymptotic analysis of the CC estimator
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, $\bf{W}$}, that minimizes the MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
We return to the example of linear regression to further break down Corollary 2.1.
Example 1 Continued (Linear Regression). For linear regression, the difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}$
while
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
In contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Returning to the example of homoskedasticity, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ always improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not. In this setting,
@@
$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}$
and
@@
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will only be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$ and the gain in efficiency will only scale in this manner with the additional condition that $\rho_{Y,\widehat{Y}}=\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$.
@@ 4 Related literature
Before evaluating the empirical properties of the CC estimator, we provide a brief overview of several closely related areas. Existing PB inference literature has noted connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error. 21, 22, 23, 32, 26, 33, 34 Here we establish additional connections between augmentation-based estimators and recently proposed PB inference approaches. We also review some modern and classical methods from the economics and statistics literature dating back to the 1960s that directly relate to PB inference. These connections underscore that aspects of this modern problem have been previously examined, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, ongoing research in PB inference would benefit from a more comprehensive review of these and other related approaches.
4.1 Extensions of PPI
We first return to the PPI++ and PSPA methods introduced in Section 2.4 that similarly aim to improve the efficiency of the PPI estimator. In contrast to the PPI and CC estimators that augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with information from $\widehat{Y}$, PPI++ and PSPA directly incorporate $\widehat{Y}$ into the estimation equation for $\bm{\beta}^{*}$. Specifically, they aim to find the solution to the augmented estimating equation
$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})-{\bf{V}}\left\{\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta})-\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta})\right\}=0$ (2)
where $\bf{V}$} is a matrix that controls the contribution of $\widehat{Y}$ into estimation of $\bm{\beta}^{*}$Similar to the PPI estimator, the authors propose to use only the unlabeled data in the augmentation term. We provide further discussion of this point in the Appendix.. For PPI++, ${\bf{V}}^{\text{\tiny{PPI++}}}=\lambda{\bf{I}}$, where $\lambda$ is selected to minimize the trace of the variance-covariance matrix of the corresponding estimator. For PSPA, ${\bf{V}}^{\text{\tiny{PSPA}}}=\text{diag}({\bf{v}})$ where $\bf{v}$} is instead selected to minimize element-wise variance. While both of these approaches fall within the AIPW framework by construction, we again return to our example of linear regression to connect these approaches to augmentation-based estimators and, in turn, show that the CC method is preferred over both PSPA and PPI++. We leave further comparisons of the three methods, particularly for general Z-estimation problems, to the Appendix and highlight some of our findings in our numerical studies in Section 5.
Example 1 Continued (Linear Regression). The PSPA and PPI++ estimators are asymptotically equivalent to augmentation-based estimators. That is,
$\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}\approx\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{PPI++}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})\quad\mbox{and}\quad\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}\approx\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{PSPA}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$
where $\widehat{{\bf{W}}}^{\text{\tiny{PPI++}}}={\bf{A}}\widehat{{\bf{V}}}^{\text{\tiny{PPI++}}}{\bf{A}}^{-1}$, $\widehat{{\bf{W}}}^{\text{\tiny{PSPA}}}={\bf{A}}\widehat{{\bf{V}}}^{\text{\tiny{PSPA}}}{\bf{A}}^{-1}$, and $\widehat{{\bf{V}}}^{\text{\tiny{PPI++}}}$ and $\widehat{{\bf{V}}}^{\text{\tiny{PSPA}}}$ are defined in the Appendix. From Corollary 2.2, the CC estimator is asymptotically at least as efficient as both PPI++ and PSPA. In fact, both can be viewed as special cases of the CC estimator: PSPA restricts $\bf{W}$} to be diagonal, while PPI++ further constrains the diagonal elements to be equal.
@@ 4.2 Seemingly unrelated regression
The augmentation-based representation not only unifies the PPI++, PSPA, and CC estimators under a common framework, but also connects to several other estimators developed for linear regression, both historically and in recent work. In particular, the framework underlies seemingly unrelated regression (SUR), a method introduced by economist Arnold Zellner in 1962 for the joint estimation of multiple related regression models. 35, 36, 37. Analogously to the current setting, the regression models have different outcome variables, but may have correlated errors. SUR leverages the correlation between the error terms to produce more efficient estimators compared to estimating each regression separately. More formally, SUR considers the regression models
$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}_{0}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}_{0}+\epsilon^{\prime}\quad$
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}_{0}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe 38 and is given by:
@@
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$, and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on $\bf{X}$}, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$. 25 Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given $\bf{X}$} is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}_{0}$ and fully efficient under the assumed parametric model. 38 In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ due to Corollary 2.2.
The SUR estimator was recently applied in a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment. 12 The proposed method, SynSurr, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al 14 proposed the POP-GWAS estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
@@ 4.3 Augmented estimation for quantifying a treatment effect
The augmented estimation approach has also been widely used for treatment effect estimation.39, 40, 41, 42, 43, 29, 44 Many of these approaches build on methods from survey sampling.45, 46 For example, Tsiatis et al 41 consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. The goal is to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information from baseline covariates $\bf{Z}$} for improving efficiency, Tsiatis et al 41 propose an augmented estimator of the treatment effect given by
@@
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of $\bf{Z}$} for $k=0,1$. As $X$ and $\bf{Z}$} are independent due to randomization, the augmentation term in 3 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. The gain in efficiency similarly arises from the potential correlation between $Y$ and $\bf{Z}$}. This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomes47 and recurrent events.31

### v6

@@ 3.2 Asymptotic analysis of the CC estimator
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g.,Bickel and Doksum,bickel2015mathematical page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, $\bf{W}$}, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Returning to the previous example, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
Example 1 continued (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ under homoskedasticity) We revisit the case of homoskedasticity and recall that
@@
$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}.$
Theorem 2 can be used to show that
@@
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will only be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$ and the gain in efficiency will only scale in this manner with the additional condition that $\rho_{Y,\widehat{Y}}=\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$.
@@ 4 Related literature
Before evaluating the empirical properties of the CC estimator, we provide a brief overview of several closely related areas. Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error. angelopoulos2023prediction, angelopoulos2023ppi++, miao2023assumption, zrnic2024active, zrnic2024cross, angelopoulos_note_2024, ji2025predictions We identify additional connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s. These connections underscore that aspects of the PB inference problem have been previously examined, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, ongoing research in PB inference would benefit from a more comprehensive review of these and other related approaches.
@@ 4.1 Seemingly unrelated regression
Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models. zellner1962efficient, swamy1975bayesian, schmidt1977estimation The regression models have different outcome variables, but are assumed to have correlated errors. SUR leverages the correlation between the error terms to produce more efficient estimators compared to estimating each regression separately. In the current setting, the SUR problem considers the regression models
$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}+\epsilon^{\prime}\quad$
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe conniffe1985estimating and is given by:
@@
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$, and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on $\bf{X}$}, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$. chen2000unified Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given $\bf{X}$} is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model. conniffe1985estimating In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
The SUR estimator was recently applied in a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment. mccaw2024synthetic The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al miao2024popgwas proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
@@ 4.2 Augmented estimation for quantifying a treatment effect
The augmented estimation approach used by the PPI and CC methods has also been widely used for treatment effect estimation.leon2003semiparametric, lu2008improving, tsiatis2008covariate, gilbert2009efficient, zhang2010increasing, tian2012covariate, yang2020combining Many of these approaches build on methods from survey sampling.cassel1976some, sarndal2003model For example, Tsiatis et al tsiatis2008covariate consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
$\mathbb{E}(Y\mid X)=\alpha+\beta X$
to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information from baseline covariates $\bf{Z}$} for improving efficiency, Tsiatis et al tsiatis2008covariate propose an augmented estimator of the treatment effect given by
@@
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of $\bf{Z}$} for $k=0,1$. As $X$ and $\bf{Z}$} are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. The gain in efficiency similarly arises from the potential correlation between $Y$ and $\bf{Z}$}. This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomeszhang2025unified and recurrent events.gronsbell2024nonparametric